Circular No. 8712

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 K1 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery observation

tabulated below) on CCD images taken with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey, noting the object to be diffuse and extended 10′′ in p.a. 240^{o}. Deeper images taken by McNaught on May 17.8 UT show a strongly condensed coma with a fan tail spanning ∼ 50^{o}. Five stacked 1-min images taken by A. C. Gilmore with the Mt. John 0.6-m f/6.4 reflector on May 18.53–18.56 show a weakly condensed coma ∼ 10′′ in diameter and a short tail suspected to the west-northwest.

2006 UT  α_{2000}  δ_{2000}  Mag.
May 17.71978  22^{h}53^{m}07.^{s}97  −60^{o}38′18′′.1  17.6

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-K13.

T = 2007 May 20.951 TT  ω = 7.^{o}358
Ω = 48.990 }2000.0
q = 1.42581 AU  i = 46.843

COMETS C/2006 B5, C/2006 B6, C/2006 C1–C3, C/2006 D2 (SOHO)

Additional Kreutz sungrazing comets have been found on SOHO website images (cf. IAUC 8694). C/2006 B5 and C/2006 C3 were described by K. Battams as tiny, stellar, and extremely faint. C/2006 B6, which was independently found by Q. Ye, was quite condensed and slightly elongated, reaching mag 4.3 on Jan. 30.638 UT at ∼ 10.8R_{⊙}. C/2006 C1 was small and stellar, reaching mag 6.4 on Feb. 8.404 at 12.4R_{⊙}. C/2006 C2 was elongated and peaked at mag ∼ 5.5, but good photometry was not possible due to a large data gap. C/2006 D2 was very small and faint in C3 images, and extremely faint and diffuse in C2 images.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 B5 Jan. 26.821 20^{h}51.^{m}6 −20^{o}08′ C3  RM 2006-J50
C/2006 B6 29.821 21 12.8 −20 29 C3 KB 2006-J50
C/2006 C1 Feb. 7.988 21 46.5 −16 41 C3  RM 2006-J50
C/2006 C2  9.513 21 58.1 −16 53 C3  SF 2006-J50
C/2006 C3  10.904 21 53.8 −15 25 C3  SF 2006-J51
C/2006 D2  16.571 22 18.6 −14 08 C3/2 HS 2006-J51
2006 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green